OpenStudy (anonymous):

Use inductive reasoning to find the 10th term in the sequence 5, 8, 11, 14, ...

OpenStudy (anonymous):

\[5,8,11,14\] Difference is 3 in all of them so the formula for the n term is simply the first term times a multiple of 3: \[ a_n=5+3n \;\;\;\;\; where \;\;n>=0.\] so the first is when n=0, namely 5, so the 10th term is when n=9, namely: \[5+3*9=5+27=32\]

OpenStudy (anonymous):

Thanks so much! You've not only helped me with the answer, but with understaning this.

OpenStudy (anonymous):

:)